OpenStudy (anonymous):

solve 3^4x-7 = 4^2x+3

OpenStudy (anonymous):

start with \[(4x-7)\ln(3)=(2x+3)\ln(4)\] and then do a raft of algebra to solve for x

OpenStudy (anonymous):

look on my profile this lady posted everything